# LongEmbed: Extending Embedding Models for Long Context Retrieval

### v2

@@ 4.2 Context Window Extension for APE-based Models
Figure 4: (Left) Arrangement of pids for extending APE-based models from 512 to 1024. (Right) Illustration of learnable (Refer to caption) and frozen (Refer to caption) position vectors when further tuning on RP / PI.
Parallel Context Windows (PCW). To process a long document with a short-context model, PCW divides the long document into multiple short chunks, processes each chunk in parallel, and aggregates their results (Ratner et al., 2023; Yen et al., 2024). In practice, we first segment $\mathcal{D}$ into chunks of $L_{o}$ tokens, then average over each chunk’s embeddings to get the embedding of $\mathcal{D}$. For simplicity, we set the overlap between adjacent chunks to 0, except for the last chunk, which conditionally overlaps with the preceding chunk to ensure it contains $L_{o}$ tokens.
Grouped Positions (GP) & Recurrent Positions (RP). Dividing inputs into chunks and processing them separately sacrifices their interaction in between. By contrast, position reorganization accommodates longer context by reusing the original position ids. To be specific, we experiment with two simple strategies: Grouped Positions and Recurrent Positions. The former groups the original position ids as such: $f_{gp}(pid)\rightarrow\lfloor pid/s\rfloor$, while the latter assigns the position ids recurrently within the range $\{0,1,...,L_{o}-1\}$, formulated as: $f_{rp}(pid)\rightarrow pid\bmod L_{o}$.
@@
4.3 Context Window Extension for RoPE-based Models
@@
Self Extend (SE). Compared with APE, RoPE operates on the query and key vectors at each layer to encode relative positions, offering enhanced flexibility for position reorganization. For each token, instead of assigning grouped relative positions to all other tokens, SelfExtend (Jin et al., 2024) re-introduces normal relative positions within the nearest neighbor window $w$, achieving improved performance. For example, consider a document of 10 tokens $\{x_{0},x_{1},...,x_{9}\}$ with a neighbor window size $w=4$ and a group size $g=2$. The relative positions for $x_{0}$ are $\{0,1,2,3,4,4,5,5,6,6\}$. For $x_{4}$, the relative positions of the other tokens are $\{-4,-3,-2,-1,0,1,2,3,4,4\}$.
NTK-Aware Interpolation (NTK). Given a scaling factor $s$, PI proportionally down-scales position index $m$ to $m/s$. In this way, the attention score $a({\bm{q}},{\bm{k}})$ defined in Equation 2 becomes $g({\bm{q}},{\bm{k}},(m-n){\bm{\theta}}/s)$. This is also equivalent to reducing the frequencies ${\bm{\theta}}$ uniformly, which may prevent the model from learning high-frequency features, as shown by the Neural Tangent Kernel (NTK) theory (Jacot et al., 2018). To remedy this, NTK-Aware interpolation (Peng & Quesnelle, 2023) scales high frequencies less and low frequencies more to spread out the interpolation pressure across multiple dimensions. This is achieved by directly altering the original $\theta_{j}=10000^{-2j/d}$ into $\theta^{\prime}_{j}=(10000\lambda)^{-2j/d}$, where $\lambda$ is conventionally chosen to be slightly greater than $s$.
@@ 5.1 Experimental Setup
Benchmarked Models. We evaluate both open-sourced and proprietary models on LongEmbed, including E5Base (Wang et al., 2022), GTEBase (Li et al., 2023), BGE-Base (Xiao et al., 2023), Contriever (Izacard et al., 2021), GTR-Base (Ni et al., 2022), E5-Mistral (Wang et al., 2023b), Jina-V2 (Günther et al., 2023), Nomic-V1 (Nussbaum et al., 2024), BGE-M3 (Chen et al., 2024), OpenAI-ada-002. For BGE-M3, we utilize dense vectors. M2 (Saad-Falcon et al., 2024) is not included in our evaluation, given its training data partly overlaps with test samples in LongEmbed.
@@ 5.2 Main Results
Table 2: Results (%) of existing and extended embedding models on LongEmbed. NQA, SFD, 2WmQA is short for NarrativeQA, SummScreenFD, 2WikiMultihopQA, respectively. We show that context window extension can effectively improve existing embedding models in handling long context.
Model Param. Synthetic (Acc@1) Real (nDCG@10) Avg.
Passkey Needle NQA QMSum SFD 2WmQA
512 Context Models
E5Base (Wang et al., 2022) 110M 38.0 28.5 25.3 23.8 74.7 55.8 41.0
E5-RoPEBase 110M 38.5 31.5 24.6 23.2 66.6 58.8 40.5
GTEBase (Li et al., 2023) 110M 31.0 24.5 28.6 21.8 55.8 47.3 34.8
BGE-Base (Xiao et al., 2023) 110M 18.0 25.3 25.6 22.4 60.3 51.7 33.9
Contriever (Izacard et al., 2021) 110M 38.5 29.0 26.7 25.5 73.5 47.3 40.1
GTR-Base (Ni et al., 2022) 110M 38.5 26.3 26.5 18.3 63.7 52.2 36.5
$\geq$ 4k Context Models
E5-Mistral (Wang et al., 2023b) 7B 71.0 48.3 44.6 43.6 96.8 82.0 64.4
Jina-V2 (Günther et al., 2023) 137M 50.3 54.5 37.9 38.9 93.5 74.0 58.2
Nomic-V1(Nussbaum et al., 2024) 137M 60.7 39.5 41.2 36.7 93.0 73.8 57.5
BGE-M3 (Chen et al., 2024) 568M 59.3 40.5 45.8 35.5 94.0 78.0 58.9
OpenAI-Ada-002 - 50.8 36.8 41.1 40.0 91.8 80.1 56.8
Our Extended Models
E5Base + Tuning (4k) 110M 67.3 41.5 30.4 35.7 95.2 69.2 56.6
E5-RoPEBase + SelfExtend (4k) 110M 73.5 53.5 32.3 39.1 91.9 74.6 60.8
E5-Mistral + NTK (32k) 7B 93.8 66.8 49.8 49.2 97.1 95.2 75.3
@@
5.3 Performance Comparison of Context Extension Methods
Figure 5: (a) Effects of different context window extension methods on E5Base and GTEBase. We found that plug-and-play methods obtain similar scores, while further tuning yields the best results. Note that for Tuning, we report results based on PI by default, as evidenced by: (b) Further ablation of tuning setups. Tuning on PI consistently outperforms RP across both models.
APE-Based Models. Figure 5(a) illustrates the impact of various context extension strategies on E5Base and GTEBase across different target context lengths. We observe that plug-and-play methods including GP, RP, LPI and PCW strategies yield comparable results with no significant disparities. On the other hand, further tuning consistently yields additional performance gains for both models, across all target context lengths. Particularly noteworthy is GTEBase, which showcases a substantial average score increase of approximately 5 points after further tuning. This suggests that freezing the original model weights and fine-tuning exclusively the added position embeddings can effectively extend the model’s context window while strictly maintaining model’s original ability.
Model Synthetic (Acc@1) Real (nDCG@10) Avg.
Passkey Needle NQA QMSum SFD 2WmQA
@@ 5.3 Performance Comparison of Context Extension Methods
Table 3: Results (%) of context window extension methods on E5-RoPEBase and E5-Mistral. NQA, SFD, 2WmQA is short for NarrativeQA, SummScreenFD, 2WikiMultihopQA, respectively.
@@ 5.4 Analysis
Tuning on PI vs. RP. Figure 5(b) compares further tuning on top of RP vs. PI. In the former approach, the initial 512 position embeddings are frozen while the remaining embeddings are tuned, whereas for the latter, the frozen / learnable embedding vectors are arranged in an interleaved manner. Our observations indicate that tuning applied to PI consistently produces superior results across both models. This superiority may be attributed to the fixed vectors acting as anchors, thereby preventing the learnable vectors from converging to suboptimal values.
Figure 6: Best results achieved by extended versions of E5 / E5-RoPE.
RoPE vs. APE. We further discuss the potential of APE / RoPE-based models for context window extension. E5Base and E5-RoPEBase are selected as the comparison subjects thanks to their shared training process, training data, and comparable performance on BEIR and LongEmbed benchmarks. At each target context length ($\{1k,2k,4k\}$), we report the best scores achieved by each model on LongEmbed, as illustrated in Figure 6. Without requiring further training, E5-RoPEBase consistently demonstrates superior performance compared to E5Base across all target lengths. Furthermore, as the target window length increases, this superiority becomes more pronounced, even surpassing the fine-tuned version of E5Base by a large margin. This suggests that RoPE-based models can better extrapolate to to longer context. Consequently, we advocate for the use of RoPE in future embedding models.
@@ Appendix B Implementation Details for Context Extension Strategies
Inference. In inference time, attention scaling (Su, 2021; Chiang & Cholak, 2022) is used by default for all tested models for better length extrapolation ability. Especially for GTEBase and E5Base tuned on PI, we use the original position ids when input length not exceeds 512. This is achived by mapping the position ids $\{0,1,...,l\}$ into $\{0,s,...,l\times s\}$, where $s$ is the scaling factor, $l<512$.

### v3

@@ 4.2 Extending APE-based Models
Figure 4: (Left) Arrangement of pids for extending APE-based models from 512 to 1,024. (Right) Illustration of learnable (Refer to caption) and frozen (Refer to caption) position vectors when further tuning on RP / PI.
Parallel Context Windows (PCW). To process a long document with a short-context model, PCW divides the long document into multiple short chunks, processes each chunk in parallel, and aggregates their results (Ratner et al., 2023; Yen et al., 2024). In practice, we first segment $\mathcal{D}$ into chunks of $L_{o}$ tokens, then average over each chunk’s embeddings to represent $\mathcal{D}$. For simplicity, we set the overlap between adjacent chunks to 0, except for the last chunk, to ensure it contains $L_{o}$ tokens.
Grouped & Recurrent Positions (GP & RP). Dividing inputs into chunks and processing them separately sacrifices their interaction in between. By contrast, position reorganization accommodates longer context by reusing the original position ids. To be specific, we experiment with two simple strategies: Grouped Positions and Recurrent Positions. The former groups the original position ids as such: $f_{gp}(pid)\rightarrow\lfloor pid/s\rfloor$, while the latter assigns the position ids recurrently, formulated as: $f_{rp}(pid)\rightarrow pid\bmod L_{o}$.
@@
4.3 Extending RoPE-based Models
@@
Self Extend (SE). Compared with APE, RoPE operates on the query and key vectors at each layer to encode relative positions, offering enhanced flexibility for position reorganization. For each token, instead of assigning grouped relative positions to all other tokens, SelfExtend (Jin et al., 2024) re-introduces normal relative positions within the nearest neighbor window $w$, achieving improved performance. For example, consider a document of 10 tokens $\{x_{0},x_{1},...,x_{9}\}$ with a neighbor window size $w=4$ and a group size $g=2$. The relative positions to $x_{0}$ are $\{0,1,2,3,4,4,5,5,6,6\}$. For $x_{4}$, the relative positions of the other tokens are $\{-4,-3,-2,-1,0,1,2,3,4,4\}$.
Model Param. CTX Len. Synthetic (Acc@1) Real (nDCG@10) Avg.
Passkey Needle NQA QMS SFD WQA
512 Context Models
E5Base (Wang et al., 2022) 110M 512 38.0 28.5 25.3 23.8 74.7 55.8 41.0
E5-RoPEBase 110M 512 38.5 31.5 24.6 23.2 66.6 58.8 40.5
GTEBase (Li et al., 2023) 110M 512 31.0 24.5 28.6 21.8 55.8 47.3 34.8
BGEBase (Xiao et al., 2023) 110M 512 18.0 25.3 25.6 22.4 60.3 51.7 33.9
Contriever (Izacard et al., 2021) 110M 512 38.5 29.0 26.7 25.5 73.5 47.3 40.1
GTRBase (Ni et al., 2022) 110M 512 38.5 26.3 26.5 18.3 63.7 52.2 36.5
$\geq$ 4k Context Models
E5-Mistral (Wang et al., 2023b) 7B 4,096 71.0 48.3 44.6 43.6 96.8 82.0 64.4
Jina-V2 (Günther et al., 2023) 137M 8,192 50.3 54.5 37.9 38.9 93.5 74.0 58.2
Nomic-V1(Nussbaum et al., 2024) 137M 8,192 60.7 39.5 41.2 36.7 93.0 73.8 57.5
BGE-M3 (Chen et al., 2024) 568M 8,192 59.3 40.5 45.8 35.5 94.0 78.0 58.9
OpenAI-Ada-002 - - 50.8 36.8 41.1 40.0 91.8 80.1 56.8
Our Extended Models
E5Base + Tuning (4k) 110M 4,096 67.3 41.5 30.4 35.7 95.2 69.2 56.6
E5-RoPEBase + SelfExtend (4k) 110M 4,096 73.5 53.5 32.3 39.1 91.9 74.6 60.8
E5-Mistral + NTK (32k) 7B 32,768 93.8 66.8 49.8 49.2 97.1 95.2 75.3
Table 2: Results (%) of existing and extended embedding models on LongEmbed. NQA, QMS, SFD, WQA is short for NarrativeQA, QMSum, SummScreenFD, 2WikiMultihopQA, respectively. We show that context window extension can effectively improve existing embedding models in processing long context.
NTK-Aware Interpolation (NTK). Given a scaling factor $s$, PI proportionally down-scales position index $m$ to $m/s$. In this way, the attention score $a({\bm{q}},{\bm{k}})$ defined in Equation 1 becomes $g({\bm{q}},{\bm{k}},(m-n){\bm{\theta}}/s)$. This is also equivalent to reducing the frequencies ${\bm{\theta}}$ uniformly, which may prevent the model from learning high-frequency features, as shown by the Neural Tangent Kernel (NTK) theory (Jacot et al., 2018). To remedy this, NTK-Aware interpolation (Peng and Quesnelle, 2023) scales high frequencies less and low frequencies more to spread out the interpolation pressure across multiple dimensions. This is achieved by directly altering the original $\theta_{j}=10000^{-2j/d}$ into $\theta^{\prime}_{j}=(10000\lambda)^{-2j/d}$, where $\lambda$ is conventionally chosen to be slightly greater than $s$.
@@ 5.1 Experimental Setup
Benchmarked Models. We evaluate both open-sourced and proprietary models on LongEmbed, including E5, GTE, BGE, Contriever, GTR, E5-Mistral, Jina-V2, Nomic-V1, BGE-M3, OpenAI-ada-002. M2 (Saad-Falcon et al., 2024) is not included in our evaluation, given its training data partly overlaps with test samples in LongEmbed.
@@ 5.2 Main Results
5.3 Comparison of Extension Methods
Figure 5: Effects of different context window extension methods on E5Base and GTEBase. We show that further tuning yields the best results.
APE-based Models. Figure 5 illustrates the impact of various context extension strategies on E5Base and GTEBase across different target context lengths. We observe that plug-and-play methods including GP, RP, PI and PCW strategies yield comparable results with no significant disparities. On the other hand, further tuning consistently yields additional performance gains for both models, across all target context lengths. Particularly noteworthy is GTEBase, which showcases a substantial average score increase of approximately 5 points after further tuning. This suggests that freezing the original model weights and fine-tuning exclusively the added position embeddings can effectively extend the model’s context window while strictly maintaining model’s original ability.
Figure 6: (a) Performance gain after tuning on PI / RP, compared with the original model. (b) Best results achieved by extended versions of E5Base / E5-RoPEBase.
Model Synthetic Real Avg.
P N NQA QMS SFD WQA
@@ 5.3 Comparison of Extension Methods
Table 3: Results (%) of context window extension methods on E5-RoPEBase and E5-Mistral. For datasets, P, N, NQA, QMS, SFD, WQA is short for Passkey, Needle, NarrativeQA, QMSum, SummScreenFD, 2WikiMultihopQA. For extension methods, PCW, GP, PI, SE, NTK are short for Parallel Context Windows, Grouped Positions, Linear Position Interpolation, SelfExtend, and NTK-Aware Interpolation, respectively.
@@ 5.4 Analysis
Tuning on PI vs. RP. Figure 6(a) compares further tuning on top of RP vs. PI. In the former approach, the initial 512 position embeddings are frozen while the remaining embeddings are tuned, whereas for the latter, the frozen / learnable embedding vectors are arranged in an interleaved manner. We observe that tuning on PI consistently produces superior results on both GTEBase and E5Base. A possible explanation is that fixed vectors in PI serve intrinsically as anchors, preventing the learnable vectors from converging to suboptimal values.
RoPE vs. APE. We further discuss the potential of APE / RoPE-based models for context window extension. E5Base and E5-RoPEBase are selected as the comparison subjects thanks to their shared training process, training data, and comparable performance on BEIR and LongEmbed benchmarks. At each target context length ($\{1k,2k,4k\}$), we report the best scores achieved by each model on LongEmbed, as illustrated in Figure 6(b). Without requiring further training, E5-RoPEBase consistently demonstrates superior performance compared to E5Base across all target lengths. Furthermore, as the target window length increases, this superiority becomes more pronounced, even surpassing the fine-tuned version of E5Base by a large margin. This suggests that RoPE-based models can better extrapolate to to longer context. Consequently, we advocate for the use of RoPE in future embedding models.
@@ 6 Conclusion
Ethics Statement
This work fully complies with the ACL Ethics Policy. We declare that there are no ethical issues in this paper, to the best of our knowledge.
Acknowledgement
We thank the anonymous reviewers for their helpful comments on this paper. We thank Xueguang Ma, Niklas Muennighoff, and Kenneth Enevoldsen for their thoughtful discussion and assistance in integrating LongEmbed into MTEB. This work was partially supported by National Natural Science Foundation of China (No. 62476010).
@@ Appendix B Implementation Details for Context Extension Strategies
Inference. In inference time, attention scaling (Su, 2021; Chiang and Cholak, 2022) is used by default for all tested models for better length extrapolation ability. Especially for GTEBase and E5Base tuned on PI, we use the original position ids when input length not exceeds 512. This is achived by mapping the position ids $\{0,1,...,l\}$ into $\{0,s,...,l\times s\}$, where $s$ is the scaling factor, $l<512$.
@@ Appendix C Further details on LongEmbed
Appendix D BM25 Results on LongEmbed
Method Synthetic Real Avg.
P N NQA QMS SFD WQA
BM25 100 95.3 71.5 81.3 97.6 96.5 90.4
E5-Mistral 71.0 48.3 44.6 43.6 96.8 82.0 64.4
+NTK (32k) 93.8 66.8 49.8 49.2 97.1 95.2 75.3
Table 7: BM25 Results on LongEmbed. P, N, NQA, QMS, SFD, WQA is short for Passkey, Needle, NarrativeQA, QMSum, SummScreenFD, 2WikiMultihopQA.
Table 7 shows the scores of BM25 on LongEmbed, along with those of the best-performing long context embedding model, E5-Mistral. The significant gap between BM25 and E5-Mistral highlights substantial room for improvement in current long context embedding models.